Circular No. 9018

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 B5 (NEAT) G. Hug, Scranton, KS, U.S.A., reports his recovery of comet P/2002

O8 (cf. IAUC 7949) on images obtained with the 0.56-m reflector at his Sandlot Observatory. The indicated correction to the prediction on MPC 59599 is ΔT = −0.34 day.

2009 UT α_{2000} δ_{2000} Mag. Observer
Jan. 22.49485 15^{h}17^{m}07.^{s}22 − 9^{o}52′16′′.5 20.3 Hug
22.51117 15 17 08.08 − 9 52 18.0 19.7  _{”}
Feb. 2.48785 15 25 27.25 −10 19 33.3 19.9  _{”}
2.51010 15 25 28.13 −10 19 35.9 20.2  _{”}

The following linked orbital elements by B. G. Marsden and an ephemeris appear on MPEC 2009-C19.

Epoch = 2010 June 13.0 TT

T = 2010 June 8.0390 TT  ω = 222.^{o}4526
e = 0.201114  Ω = 75.4410 }2000.0
q = 3.213378 AU  i = 12.7900

a = 4.022322 AU n^{o} = 0.1221768 P = 8.07 years

COMETS C/2008 T6–T10 (SOHO) Further to IAUC 9014, additional near-sun presumed comets have

been found on SOHO website images; all were Kreutz sungrazers except for C/2008 T6 (Meyer group) and C/2008 T7 (no known group). Those objects visible in C3 images appeared stellar therein. C/2008 T6 was small and faded quickly (peak mag ∼ 7). C/2008 T7 was tiny (mag ∼ 7.5–8). C/2008 T8 peaked at mag ∼ 6.5–7 in C3 images, but was quite diffuse with a short, diffuse tail in C2 images. C/2008 T9 and C/2008 T10 were of mag 7.5–8 in C3 images; the former was diffuse and very faint in C2 images, while the latter was very diffuse and slightly elongated in C2 images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 T6 Oct. 2.379 12^{h}31.^{m}1 − 2^{o}17′ C2  BZ 2009-A59
C/2008 T7 3.868 12 42.4 − 2 46 C2 RK 2009-A59
C/2008 T8  5.279 12 26.6 − 6 13 C3/2 JR 2009-A59
C/2008 T9  5.304 12 31.8 − 6 02 C3/2 MK 2009-A59
C/2008 T10  11.846 12 57.8 − 8 32 C3/2 TH 2009-A59
2009 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green